S/D/E Observing - LUCI2 - UT20151124

### LUCI2 "Service" Observing - UT 20151124
luci2.20151124.NNNN.fits

Summary:
  • INAF: Antonniucci_7 v2492_cyg Completed, significant edits needed
  • LBTB:Wuyts UDSK VClose to moon, skipped
  • LBTO:Rothberg NGC_34 Completed, use shorter acq times for bright targets
  • INAF/Iovino 2015_2016_5_zJ_spec Problem w/LMS file, skipped
  • LBTO/DThompson hero005 Completed
  • LBTO:Rothberg UGC_2238 Completed, use shorter acq times for bright targets
  • INAF/Antonucci v1647_ori Completed, significant edits needed
  • LBTB:Wuyts M2333cosk2 Did 1.83 hrs (22 x 300s)
Problems:
  • Had to restart the scheduler twice. It might be related to inserting/skipping pauses. Note sent to Alexander
  • Mask with only 1 offset star (RTD requires boxes around all alignment stars for now), won't work
  • Several scripts heavily saturated the targets...use the ETC!
  • LUCI2 G210 grating having trouble stabilizing for K-band CWL < 2.24um. Note added to "writing scripts" page
0001 0002 1x2.74s dark to check readnoise (LS0.5" in FPU)

12:34 Init all units in LUCI2 ... OK

Mirror is 2.6 degrees warmer than outside
Clear at sunset

### INAF: Antonniucci_7
README says 0.75" slit, scripts made up with 1.00 arcsec. Bright sources, so I
switched to LS_0.50arcsec as they can always degrade the resolution.

12:50 v2492_cyg_tell_LS0.5.xml
  • Initial seeing on the guider is 0.6"
  • zJspec is bright, HKspec likely saturated in 5.0s
01:03 v2492_cyg_tell_LS0.5_3s.xml
  • Modified to do DIT=2.75s
  • Image quality on the acq image is terrible!
  • Last file: 0016
01:12 v2492_cyg_LS0.5.xml
  • Acq: +0.15" dX done manually to center on slit
  • Acq: +0.15" dX done manually again
  • 0.45" on guider images
  • zJspec data look good
  • HKspec data look close to saturation
Remade HKspec script: v2492_cyg_LS0.5_180s_HKspec.xml
  • Changed to 6x[1x180s] exposures, first offset to recenter
  • Still too close to saturation! Aborted
Remade HKspec script: v2492_cyg_LS0.5_60s_HKspec.xml
  • Changed to 6x[1x60s] exposures, first offset to recenter
  • Still too close to saturation! Aborted
Remade HKspec script: v2492_cyg_LS0.5_20s_HKspec.xml
  • Changed to 6x[3x20s] LIR/INT exposures, first offset to recenter
  • OK now, ~22000 ADU peak in the spectrum in 20s
02:33 v2492_cyg_img.xml
  • DIQ is not great
  • Completely saturated!
02:44 v2492_cyg_FeII_img.xml
  • DIQ is not great
  • 26k ADU in 2.75s
--> mv v2492_cyg* DONE

### LBTB: UDSK

02:50 M2337udsk_tell.xml
  • No README with scripts
  • No FC with slits identified for telluric star (was understandable from offsets)
  • Very close to the MOON !! Skipping this one tonight
  • Guider did not find the guide star, background very high
### Rothberg: NGC_34

02:54 HD_219833_Telluric_2pt27.xml
  • Script out of sync with instrument configuration (udsk mask still in the FPU)
  • Scheduler restarted
  • Reset GEIRS on READ2 gui (take test exposure...OK)
  • Refresh on LUCI2 gui...OK
03:02 HD_219833_Telluric_2pt27.xml re-loaded and re-sent
  • Mask moving now
  • EL~44 setting
  • 20x3.0s in K for acquisition of a telluric ???
03:17 NGC_34_V2_2pt27.xml
  • 20x3.0s for acquisition was overkill, sky also probably not needed
  • ~16k ADU peak in the blue end of the spectrum
### INAF/Iovino

04:26 2015_2016_5_zJ_spec_tell.xml
  • EL~50, rising. Guider around 1" now
  • Incorrect mask ID in script. Was ID993628, should be 993628.XXLrev
04:32 Running Corrected_zJ_spec_tell.xml
  • LMS file not in the working directory - downloaded from luci.luci
  • MOS acquisition is not working, sky image seems to be missing
04:50 Corrected_zJ_spec_tell.xml (again)
  • Cannot acquire mask. not sure where the problem is. Skipping
  • Looks like RTD only sees a single reference star for alignment
### LBTO/DThompson

05:04 hero005.xml
  • Very near moon, but the guide star was found
  • Seeing 0.8"
  • Can faintly see the target continuum in the K band
### LBTO:Rothberg

~07:00 observed telluric HD18571_Tell_2pt27


07:15 acquiring UGC_2238_2pt27
  • Seeing 0.7" on guider
  • Thin cirrus coming in from West
07:25 started spectroscopic observation

### INAF/Antonucci_7

08:30 v1647_ori_60s.xml
  • Source is 10th mag in K, 14th in J
  • Changed the spectroscopic exposures to 60s
  • 0.62" seeing
  • 7k ADU max in zJspec, saturated in HKspec
### INAF/Antonniucci_7

09:02 v1647_ori_10s_HKspec.xml
  • Original script saturates in HKspec
  • Changed to 6x10s LIR/INT HKspec only to redo it
  • 24k ADU per DIT
09:12 v1647_ori_tell_2.75s.xml
  • 29k ADU max in HKspec spectrum
09:24 v1647_ori_img_2.75s.xml
  • Guide probe is visible on detector, left of science target
  • Seeing got worse, now ~1.5"
### LBTB/Wuyts

09:33 M2333cosk2.xml
  • Could not do the telluric before...EL=17!
  • MOS acq, then +0.15" manual dX offset (no image taken)
  • G210 @ 2.20um stabilized OK
  • Seeing a bit variable, 1.0-1.3", 1.1" avg.
  • Emission-line source seen in bottom slit in all exposures I checked
12:01 M2333cosk2_tell.xml
  • Aborted...scheduler needs restart after a pause that gets reset?
  • Dithers between 3 slitlets worked well
12:13 Closing dome

### Calibrations

12:27 M2333cosk2_cal.xml
  • Starts at file 0179
  • Grating not stabilizing at 2.20um. It did not have any trouble on sky.
  • Tried manually setting various wavelengths, no luck @ 2.2um
12:49 Cal_0pt5_210K_V3_2pt27.xml
  • Failed to lock @ 2.27um
  • Accidentally moved COSK2 mask in when trying manually toi tilt grating
01:01 M2333cosk2_cal.xml
  • Failed to lock @ 2.2um
  • Continued script (log says grating unlocked)
  • Flatfields might still be OK
  • Still failing..."continue" on script continues unlocked
01:19 Cal_0pt5_210K_V3_2pt27.xml
  • Trying again...worked?
01:50 Doing some calibrations by hand

G200 LS_0.50arcsec 1.175 zJspec
  • 0232 0236 1x3.0s Halo2 lamp off
  • 0237 0241 1x3.0s Halo2 lamp on ~2.5k ADU
G200 LS_0.50arcsec 1.930 HKspec
  • 0242 0246 1x3.0s Halo2 lamp off
  • 0247 0251 1x3.0s Halo2 lamp on ~11k ADU
G200 LS_1.00arcsec 1.930
  • 0252 0256 1x30.0s Halo3 lamp off ~3.5k ADU
  • 0257 0261 1x30.0s Halo3 lamp on ~7.5k ADU
G200 LS_1.00arcsec 1.175
  • 0262 0266 1x30.0s Halo3 lamp off ~0k ADU
  • 0267 0271 1x30.0s Halo3 lamp on ~1.5kk ADU
  • 0272 0276 1x3.0s Halo2 lamp off ~0k ADU
  • 0277 0281 1x3.0s Halo2 lamp on ~4.5kk ADU
Still need to run calibrations for
Arcs, if you don't want to use OH lines



>>> Need to adjust the LUCI2 field stop before tomorrow





Topic revision: r2 - 24 Nov 2015, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback